Comment pquest SNF 2009-06-16 15:21:35: Update

eenriquez at snf.stanford.edu eenriquez at snf.stanford.edu
Tue Jun 16 15:21:36 PDT 2009


Adjusted the dark space shield height and centered the clamp.  Ran
the Oxygen recipe.  The bias voltage = -45 v,  He cooling flow= 2.3 sccm
reflected power= 0 W.
The bad news:  
   The clamp mid position is inconsistent.  We believe this is caused by the sensor block guide rod having too much slop.  Need to tighten up the guide rod.
 The load arm has disengaged from its drive mechanism. Need to reattach the load arm.   




More information about the pquest-pcs mailing list